Write the equation of the line in standard form.Use integer coefficients
y=-3/4x-1/2

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
y = -(3/4)x - 1/2
-(3/4)x - y = 1/2
( -(3/4)x - y = 1/2 ) * 4
---
answer:
-3x - 4y = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

Write the equation of the line in standard form.Use integer coefficients
y=-3/4x-1/2
-------
Multiply thru by 4 to get:
4y = -3x - 2
-----
Standard Form::
3x + 4y = -2
====================
Cheers,
Stan H.
=======================